# Differentiable Grouped Feedback Delay Networks for Learning Coupled Volume Acoustics

### v1

@@ Abstract
Rendering dynamic reverberation in a complicated acoustic space for moving sources and listeners is challenging but crucial for enhancing user immersion in extended-reality (XR) applications. Capturing spatially varying room impulse responses (RIRs) is costly and often impractical. Moreover, dynamic convolution with measured RIRs is computationally expensive with high memory demands, typically not available on wearable computing devices. Grouped Feedback Delay Networks (GFDNs), on the other hand, allow efficient rendering of coupled room acoustics. However, its parameters need to be tuned to match the reverberation profile of a coupled space. In this work, we propose the concept of Differentiable GFDNs (DiffGFDNs), which have tunable parameters that are optimised to match the late reverberation profile of a set of RIRs captured from a space that exhibits multi-slope decay. Once trained on a finite set of measurements, the DiffGFDN generalises to unmeasured locations in the space. We propose a parallel processing pipeline that has multiple DiffGFDNs with frequency-independent parameters processing each octave band. The parameters of the DiffGFDN can be updated rapidly during inferencing as sources and listeners move. We evaluate the proposed architecture against the Common Slopes (CS) model on a dataset of RIRs for three coupled rooms. The proposed architecture generates multi-slope late reverberation with low memory and computational requirements, achieving better energy decay relief (EDR) error and slightly worse octave-band energy decay curve (EDC) errors compared to the CS model. Furthermore, DiffGFDN requires an order of magnitude fewer floating-point operations per sample than the CS renderer.
Index Terms:
3D sound, acoustic signal processing, augmented reality, machine learning, optimization, reverberation.
@@ I Introduction
Rendering dynamic reverberation for moving sources and listeners in complex environments is a challenging problem, but essential for immersive extended reality (XR) applications [1, 2]. In simple shoebox-shaped rooms with uniform absorption, late reverberation is typically approximated by an isotropic and diffuse sound field [3, 4], and can be well-modeled using filtered noise tails or tuned feedback delay networks (FDNs) [5, 6]. However, in coupled spaces with different reverberation characteristics and rooms with non-uniform absorption, late reverberation becomes inhomogeneous and anisotropic, giving rise to effects such as multi-slope decay [7, 8]. Acoustic renderers in Augmented Reality (AR) applications need to model these spatially varying dynamics from limited available data, whilst being efficient and lightweight [9].
@@
The Common Slopes (CS) model [15] offers a data-driven alternative for modelling late reverberation in coupled spaces. It represents multi-slope decays using a small set of shared decay rates and spatially varying amplitudes, enabling compact RIR modeling. During real-time rendering, position-dependent equalisers adjust the amplitudes, while static modal filters reproduce the decays [16]. Although effective, this approach relies on computationally expensive modal reverberators, requiring hundreds of modes per band, and cannot generalise to unmeasured positions. We recently proposed interpolating CS amplitudes to unseen positions using multilayer perceptrons (MLPs) [17], but synthesising reverberation still requires time-varying convolution.
Scerbo et al. [18] leverage the CS model to simplify multi-slope reverberation rendering in a virtual acoustic space using an FDN. The method uses acoustic radiance transfer (ART) to derive modal representations in an FDN (ART-FDN) [19], where the common decay rates are encoded in real poles, and listener/source effects are modelled via filters. Similarly, the Coupled Volume Scattering Delay Network [20] targets coupled shoebox-shaped spaces. These approaches are promising for virtual environments but not directly applicable to real measured spaces.
Recent advances in differentiable FDNs have shown that network parameters can be learned directly from measured RIRs using gradient descent. Approaches include perceptually-informed loss functions [21, 22, 23], end-to-end training for target reverberation characteristics [24], and multichannel MIMO architectures trained on spatial datasets [25]. However, most of these models are agnostic to spatial coordinates, limiting their ability to generalise to unseen positions.
@@
The rest of this paper is organised as follows. In Sec. II, we explain the architecture and parameters of the GFDN. In Sec. III, we derive the RIRs synthesised by the GFDN, and compare those with those of the CS model. In Sec. IV, we discuss the configuration of the learnable parameters of the DiffGFDN and the loss functions minimised to optimise them. In Sec. V, we propose a band-wise processing architecture with multiple frequency-independent DiffGFDNs, and derive its computational and memory requirements. In Sec. VI, we evaluate the proposed subband architecture on a three-coupled room dataset, and compare the energy decay curve (EDC) and energy decay relief (EDR) errors with those of the CS model for different amounts of training data. Sec. VII concludes the paper.
@@ II Architecture
Figure 1: Differentiable Grouped FDN architecture for $G=3$ groups. The bold lines represent multichannel signals, and the different colours represent different groups. $\mathbf{s}_{p}$ and $\mathbf{r}_{q}$ represent source and receiver positions, respectively, and $H_{pq}(z)$ and $\hat{H}_{pq}(z)$ are the reference and predicted transfer functions, respectively, for the position pair ($\mathbf{s}_{p},\mathbf{r}_{q}$).
@@
$\bm{b}(\mathbf{x},z)$ $=\bm{b}\odot\left(\bm{g}_{i}(\mathbf{x},z)\otimes\bm{1}\right),\quad\bm{c}(\mathbf{x},z)=\bm{c}\odot\left(\bm{g}_{o}(\mathbf{x},z)\otimes\bm{1}\right),$
@@
      $\Phi\Phi^{H}=\mathbf{I},\quad\mathbf{M}_{k}^{H}\mathbf{M}_{k}=\mathbf{I},$
@@ III-A RIRs Synthesised by the Common Slopes Model
To get the reverberated output in real-time, a modal reverberator generates $M_{b}$ modes in each band, with a centre frequency of $f_{b}$ Hz. The mode frequencies, $f_{m,b}$, are logarithmically distributed around the centre frequency, such that, $\frac{f_{b}}{\sqrt{2}}<f_{m,b}<\sqrt{2}f_{b}$. The mode phases, $\varphi_{m,b}$, are randomly generated between $0,2\pi]$ radians.
@@
$\mathcal{A}_{k}(\mathbf{x},f)=\sqrt{A_{k,b}(\mathbf{x})}\quad\text{ for }\frac{f_{b}}{\sqrt{2}}<f<\sqrt{2}f.$
@@ III-B Structure of the Feedback Matrix
A feedback matrix with no coupling between groups allows deriving close relationships between the proposed architecture and the CS model. We started with a no-coupling initialisation (diagonal coupling matrix), and post-training, the learned coupling matrix remained diagonal. When we tried to enforce coupling amongst the groups the EDC fit was poor.
@@
We now derive the transfer function of the GFDN for a block-diagonal feedback matrix with no coupling. Let the input-output gains for the $k^{\text{th}}$ FDN in the GFDN be denoted as $\bm{b}_{k},\ \bm{c}_{k}\in\mathbb{R}^{N^{\prime}\times 1}$ and let the feedback matrix be $\mathbf{M}_{k}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$. The delay line lengths are denoted by $\bm{m}_{k}\in\mathbb{Z}^{+N^{\prime}\times 1}$ and the diagonal matrix formed from the delays is $\bm{D_{m_{k}}}(z)=\text{diag}\left(z^{-\bm{m_{k}}}\right)$. The absorption matrix is, $\bm{\Gamma}_{k}(z)=\text{diag}(\gamma_{k}(z)^{\bm{m}_{k}})$ The transfer function of the GFDN with the block diagonal feedback matrix ($\bm{\Phi}=\bm{I}_{G}$) can be written as
@@
$\bm{b}_{k}(\mathbf{x},z)$ $=\bm{b}_{k}g_{i,k}(\mathbf{x},z),\qquad\bm{c}_{k}(\mathbf{x},z)=\bm{c}_{k}g_{o,k}(\mathbf{x},z).$
@@ III-C RIRs Synthesised by Decoupled GFDN
Here, $\lambda_{m,k,b}$ are the poles, and $\rho_{m,k,b}$ are the residues for the $m^{\text{th}}$ mode in the $k^{\text{th}}$ group and $b^{\text{th}}$ frequency band. The pole frequencies follow, $\frac{f_{b}}{\sqrt{2}}<\frac{\angle{\lambda_{{m,k,b}}}f_{s}}{2\pi}<\sqrt{2}f_{b}$. In Appendix A, we simplify this expression further and show that the pole magnitudes are approximately constant over a frequency band, i.e., , $|\lambda_{m,k,b}|\approx|\lambda_{k,b}|=\exp\left(\frac{-6.91}{f_{s}T_{60_{k,b}}}\right)$. Hence, we can write the impulse response as
@@ IV Proposed Parameter Learning
Starting from a set of RIRs measured at different source-receiver positions, we would like to train the parameters of the GFDN, so that it can generalise late reverberation at unseen source-receiver positions. In this section, we describe the fixed and learnable parameters of the GFDN, and the differentiable loss functions used for optimisation via stochastic gradient descent. We refer to the GFDN with learnable parameters as the DiffGFDN. The loss functions are designed to capture the energy decay profile of the late reverberation being modelled.
@@ IV-A Fixed Parameters
  • The absorption filters, $\bm{\gamma}(z)$, are derived from the common decay times in octave bands. The common decay times can be found by the method suggested in [15], where the individual RIRs’ reverberation times in octave bands are first found using a neural network [32], and then clustered into $G$ groups. The common decay times are converted to absorption gains using (2). The absorption filters are modelled with a graphic equaliser (GEQ) with one band per octave, and optimised command gains. The GEQ is designed as a cascade of biquads filters with fixed center frequencies at the octaves bands. The optimum command gains are found via a constrained linear least squares method [26], by minimising the $\ell_{2}$-norm of the deviation of the GEQ’s magnitude response of the desired absorption filter gains, $\bm{\gamma}_{k,b_{\text{dB}}}(\omega_{\mathrm{P}})=\frac{-60\bm{m}}{f_{s}T_{60}(\omega_{\mathrm{P}})}$, where $T_{60}(\omega_{\mathrm{P}})$ are the reverberation times at logarithmically placed control frequencies, $\omega_{\mathrm{P}}$. For more information on the structure of a GEQ, please see [33].
@@ IV-B Learnable Parameters
    $\bm{A}$ $=\begin{bmatrix}\mathbf{M}_{1}^{2}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots\\ \bm{0}&\cdots&\mathbf{M}^{2}_{G}\end{bmatrix}$, (14)
@@
    where $\mathbf{W}_{k,\text{Tr}}$ is the upper triangular part of a unconstrained real-valued matrix, $\mathbf{W}_{k}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$ and $\exp(.)$ denotes the matrix exponential. The matrices, $\mathbf{W}_{k}$, are learned via gradient descent.
@@
    where the sine and cosine functions are applied element-wise to $\mathbf{x}$, and $N_{f}$ is the number of input frequencies. The output of the MLP are either scalar gains, or state variable filter (SVF) coefficients that are converted to parametric equalisers (PEQ) [24] that represent the source and receiver filters, $\bm{g}_{i}(\mathbf{x},z),\bm{g}_{o}(\mathbf{x},z)$. More details on the MLP output are provided in Sec. V.
@@ IV-C1 Energy Decay Loss Functions
To match the energy decay of a measured transfer function, $H_{pq}(z)$, at a source-receiver location, $(\mathbf{s}_{p},\mathbf{r}_{q})$, we minimise the normalised energy decay relief (EDR) loss between the DiffGFDN’s output and the reference RIR at each location [35]. The EDR characterises the energy decay of an RIR as a function of frequency and time, and is defined by the squared magnitude of the short-time Fourier transform (STFT) of the impulse response [36]. We form an EDR loss function, $\mathcal{L}_{\text{EDR}_{pq}}$, based on the EDR of the reference transfer function, $\text{EDR}_{pq}(k,n)$, and the EDR obtained from the GFDN’s transfer function, $\widehat{\text{EDR}}_{pq}(k,j)$, where $j$ is the time index and $k$ is the frequency bin of the STFT which has a total of $J$ time indices.
@@
$\mathcal{L}_{\text{EDC}_{pq}}$ $=\frac{1}{Q}\sum_{t=1}^{Q}\left|\text{EDC}_{pq}(n)-\widehat{\text{EDC}}_{pq}(n)\right|.$

### v2

@@ Abstract
Rendering dynamic reverberation in a complicated acoustic space for moving sources and listeners is challenging but crucial for enhancing user immersion in extended-reality (XR) applications. Capturing spatially varying room impulse responses (RIRs) is costly and often impractical. Moreover, dynamic convolution with measured RIRs is computationally expensive with high memory demands, typically not available on wearable computing devices. Grouped Feedback Delay Networks (GFDNs), on the other hand, allow efficient rendering of coupled room acoustics. However, its parameters need to be tuned to match the reverberation profile of a coupled space. In this work, we propose the concept of Differentiable GFDNs (DiffGFDNs), which have tunable parameters that are optimised to match the late reverberation profile of a set of RIRs captured from a space that exhibits multi-slope decay. Once trained on a finite set of measurements, the DiffGFDN interpolates to unmeasured locations in the space. We propose a parallel processing pipeline that has multiple DiffGFDNs with frequency-independent parameters processing each octave band. The parameters of the DiffGFDN can be updated rapidly during inferencing as sources and listeners move. We evaluate the proposed architecture against the Common Slopes (CS) model on a dataset of RIRs for three coupled rooms. The proposed architecture generates multi-slope late reverberation with low memory and computational requirements, achieving a better energy decay relief (EDR) error and slightly worse octave-band energy decay curve (EDC) errors compared to the CS model. Furthermore, DiffGFDN requires an order of magnitude fewer floating-point operations per sample than the CS renderer.
@@ I Introduction
Rendering dynamic reverberation for moving sources and listeners in complex environments is a challenging problem, but essential for immersive extended reality (XR) applications [42, 43]. In simple cuboid rooms with uniform absorption, late reverberation is typically approximated by an isotropic and diffuse sound field [33, 27], and can be well-modeled using filtered noise tails or tuned feedback delay networks (FDNs) [53, 44]. However, in coupled spaces with different reverberation characteristics and rooms with non-uniform absorption, late reverberation becomes inhomogeneous and anisotropic, giving rise to effects such as multi-slope decay [3, 57]. Acoustic renderers in Augmented Reality (AR) applications need to model these spatially varying dynamics from limited available data, whilst being efficient and lightweight [58].
@@
The Common Slopes (CS) model [22] offers a data-driven alternative for modelling late reverberation in coupled spaces. It represents multi-slope decays using a small set of shared decay rates and spatially varying amplitudes, enabling compact RIR modeling. During real-time rendering, position-dependent equalisers adjust the amplitudes, while static modal filters reproduce the decays [21]. Although effective, this approach relies on computationally expensive modal reverberators, requiring hundreds of modes per band, and cannot extend to unmeasured positions. We recently proposed interpolating CS amplitudes to unseen positions using multilayer perceptrons (MLPs) [16], synthesising the late tail with filtered noise shaped by the CS parameters. However, rendering still requires time-varying convolution with the synthesised tail. In contrast, the method proposed in this paper also learns its parameters from discrete RIR measurements but eliminates convolution by replacing it with a recursive delay network.
Scerbo et al. leverage the CS model to simplify multi-slope reverberation rendering in a virtual acoustic space using FDNs [46, 47]. The method uses acoustic radiance transfer (ART) to derive modal representations in an FDN (ART-FDN) [7], where the common decay rates are encoded in real poles, and listener/source effects are modelled via dynamic gains. Smaller FDNs with fewer delay lines and different weightings are used to render late reverberation in real-time for multiple dynamic sources and listeners. Similarly, the Coupled Volume Scattering Delay Network [5] targets coupled shoebox-shaped spaces. In [32], the authors propose an FDN-based, geometrically inspired reverberator for rendering coupled volume acoustics. Each sub-room in the coupled space is represented by an independent FDN, and the overall reverberation is obtained by summing their outputs, similar to the GFDN. Coupling between the volumes is modelled through weighting coefficients applied to each FDN as well as cross-mixing between their outputs. Directional late reverberation is captured by placing spatially warped virtual reverberation sources around the listener. These approaches are promising for virtual environments but not directly applicable to real measured spaces.
Recent advances in differentiable FDNs have shown that network parameters can be learned directly from measured RIRs using gradient descent. Approaches include perceptually-informed loss functions [12, 10, 40], end-to-end training for target reverberation characteristics [35], and multichannel MIMO architectures trained on spatial datasets [20]. However, most of these models are agnostic to spatial coordinates, limiting their ability to interpolate to unseen positions. Recently proposed differentiable Scattering Delay Networks [41] can predict RIRs at unseen positions, however, they still require approximate knowledge of the room geometry, work only for shoebox rooms and cannot yet model multi-slope late reverberation observed in coupled spaces.
@@
The rest of this paper is organised as follows. In Sec. II, we explain the architecture and parameters of the GFDN. In Sec. III, we derive the RIRs synthesised by the GFDN, and compare those with those of the CS model. In Sec. IV, we discuss the configuration of the learnable parameters of the DiffGFDN and the loss functions minimised to optimise them. In Sec. V, we propose a band-wise processing architecture with multiple frequency-independent DiffGFDNs, and derive its computational and memory requirements. In Sec. VI, we evaluate the proposed subband architecture on a three-coupled room dataset with an ablation study, and compare the energy decay curve (EDC) and energy decay relief (EDR) errors with those of the CS model for different amounts of training data. In Sec. VII, we discuss the benefits and limitations of the proposed method for acoustic rendering in AR. Sec. VIII concludes the paper.
@@ II Architecture
Figure 1: Differentiable Grouped FDN architecture for $G=3$ groups. The bold lines represent multichannel signals, and the different colours represent different groups, $\mathbf{s}_{p}$ and $\mathbf{r}_{q}$ represent source and receiver positions, respectively, and $H_{pq}(z)$ and $\hat{H}_{pq}(z)$ are the reference and predicted transfer functions, respectively, for the position pair ($\mathbf{s}_{p},\mathbf{r}_{q}$).
@@
$\bm{b}(\mathbf{x},z)$ $=\bm{b}\odot\left(\bm{g}_{i}(\mathbf{x},z)\otimes\bm{1}\right),$
$\bm{c}(\mathbf{x},z)$ $=\bm{c}\odot\left(\bm{g}_{o}(\mathbf{x},z)\otimes\bm{1}\right),$
@@
      $\Phi\Phi^{H}=\mathbf{I},\qquad\mathbf{M}_{k}^{H}\mathbf{M}_{k}=\mathbf{I},$
@@ III-A RIRs Synthesised by the Common Slopes Model
To get the reverberated output in real-time, a modal reverberator generates $M_{b}$ modes in each band, with a centre frequency of $f_{b}$ Hz. In total, there are $B$ frequency-bands. The mode frequencies, $f_{m,b}$, are logarithmically distributed around the centre frequency, such that, $\frac{f_{b}}{\sqrt{2}}<f_{m,b}<\sqrt{2}f_{b}$. The mode phases, $\varphi_{m,b}$, are randomly generated between $0,2\pi]$ radians.
@@
$\mathcal{A}_{k}(\mathbf{x},f)=\sqrt{A_{k,b}(\mathbf{x})}\quad\text{ for }\frac{f_{b}}{\sqrt{2}}<f<\sqrt{2}f_{b}.$
@@ III-B Structure of the Feedback Matrix
A feedback matrix with no coupling between groups allows deriving close relationships between the proposed architecture and the CS model.
@@
We now derive the transfer function of the GFDN for a block-diagonal feedback matrix with no coupling. Let the input-output gains for the $k^{\text{th}}$ FDN in the GFDN be denoted as $\bm{b}_{k},\ \bm{c}_{k}\in\mathbb{R}^{N^{\prime}\times 1}$ and let the feedback matrix be $\mathbf{M}_{k}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$. The delay line lengths are denoted by $\bm{m}_{k}\in\mathbb{Z}^{+N^{\prime}\times 1}$ and the diagonal matrix formed from the delays is $\bm{D_{m_{k}}}(z)=\text{diag}\left(z^{-\bm{m_{k}}}\right)$. The absorption matrix is $\bm{\Gamma}_{k}(z)=\text{diag}(\gamma_{k}(z)^{\bm{m}_{k}})$. The transfer function of the GFDN with the block diagonal feedback matrix ($\bm{\Phi}=\bm{I}_{G}$) can be written as
@@
where
$\bm{b}_{k}(\mathbf{x},z)=\bm{b}_{k}g_{i,k}(\mathbf{x},z)~,~\bm{c}_{k}(\mathbf{x},z)=\bm{c}_{k}g_{o,k}(\mathbf{x},z).$
@@ III-C RIRs Synthesised by Decoupled GFDN
Here, $\lambda_{m,k,b}$ are the poles, and $\rho_{m,k,b}$ are the residues for the $m^{\text{th}}$ mode in the $k^{\text{th}}$ group and $b^{\text{th}}$ frequency band. The pole frequencies follow
$\frac{f_{b}}{\sqrt{2}}<\frac{\angle{\lambda_{{m,k,b}}}f_{s}}{2\pi}<\sqrt{2}f_{b}~.$
In Appendix A, we simplify this expression further and show that the pole magnitudes are approximately constant over a frequency band, i.e., , $|\lambda_{m,k,b}|\approx|\lambda_{k,b}|=\exp\left(\frac{-6.91}{f_{s}T_{60_{k,b}}}\right)$. Hence, we can write the impulse response as
@@ IV Proposed Parameter Learning
Starting from a set of RIRs measured at different source-receiver positions, we would like to train the parameters of the GFDN, so that it can interpolate late reverberation at unseen source-receiver positions. In this section, we describe the fixed and learnable parameters of the GFDN, and the differentiable loss functions used for optimisation via stochastic gradient descent. We refer to the GFDN with learnable parameters as the DiffGFDN. The loss functions are designed to capture the energy decay profile of the late reverberation being modelled.
@@ IV-A Fixed Parameters
  • The absorption filters, $\bm{\gamma}(z)$, are derived from the common decay times in octave bands. The common decay times can be found by the method suggested in [22], where the individual RIRs’ reverberation times in octave bands are first found using a neural network [Götz_Falcón_Pérez_Schlecht_Pulkki_2022], and then clustered into $G$ groups. The common decay times are converted to absorption gains using (2). The absorption filters are modelled with a graphic equaliser (GEQ) with one band per octave, and optimised command gains. The GEQ is designed as a cascade of biquad filters with fixed center frequencies at the octave bands. The optimum command gains are found via a constrained linear least squares method [48], by minimising the $\ell_{2}$-norm of the deviation of the GEQ’s magnitude response of the desired absorption filter gains,
    $\bm{\gamma}_{k,b_{\text{dB}}}(\omega_{\mathrm{P}})=\frac{-60\bm{m}}{f_{s}T_{60}(\omega_{\mathrm{P}})}~,$
    where $T_{60}(\omega_{\mathrm{P}})$ are the reverberation times at logarithmically placed control frequencies, $\omega_{\mathrm{P}}$. For more information on the structure of a GEQ, please see [56].
@@ IV-B Learnable Parameters
    $\bm{A}$ $=\begin{bmatrix}\mathbf{M}_{1}^{2}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots\\ \bm{0}&\cdots&\mathbf{M}^{2}_{G}\end{bmatrix}~$. (14)
@@
    where $\mathbf{W}_{k,\text{Tr}}$ is the upper triangular part of a unconstrained real-valued matrix, $\mathbf{W}_{k}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$ and $\exp(.)$ denotes the matrix exponential. The matrices, $\mathbf{W}_{k}$, are learned via gradient descent. Note that a position-invariant feedback matrix is needed for analogy with the CS model. It ensures the inner loop of the GFDN, (10), produces noise shaped by $G$ decay kernels with $G$ distinct decay profiles, which are weighted by the position-dependent source-receiver filters.
@@
    where the sine and cosine functions are applied element-wise to $\mathbf{x}$, and $N_{f}$ is the number of input frequencies. The output of the MLP are either scalar gains, or state variable filter (SVF) coefficients that are converted to parametric equalisers (PEQ) [35] that represent the source and receiver filters, $\bm{g}_{i}(\mathbf{x},z),\bm{g}_{o}(\mathbf{x},z)$. More detail on the MLP output is provided in Sec. V.
@@ IV-C1 Energy Decay Loss Functions
To match the energy decay of a measured transfer function, $H_{pq}(z)$, at a source-receiver location, $(\mathbf{s}_{p},\mathbf{r}_{q})$, we minimise the normalised energy decay relief (EDR) loss between the DiffGFDN’s output and the reference RIR at each location [39]. The EDR characterises the energy decay of an RIR as a function of frequency and time, and is defined by the squared magnitude of the short-time Fourier transform (STFT) of the impulse response [30]. We form an EDR loss function, $\mathcal{L}_{\text{EDR}_{pq}}$, based on the EDR of the reference transfer function, $\text{EDR}_{pq}(k,n)$, and the EDR obtained from the GFDN’s transfer function, $\widehat{\text{EDR}}_{pq}(k,j)$, where $j$ is the time index and $k$ is the frequency bin of the STFT which has a total of $J$ time indices:
@@
$\mathcal{L}_{\text{EDC}_{pq}}$ $=\frac{1}{Q}\sum_{n=1}^{Q}\left|\text{EDC}_{pq}(n)-\widehat{\text{EDC}}_{pq}(n)\right|.$